Hi,
I would like to calculate values based on the following given:
23 points is 100%
16 points is 60%
So by that logic, what is the percentage of 17, 18, 19, 20, 21 and 22 points?
Normally, 16 points would equal 70%, and if you assume that 16 points is 60%, then 100% would normally be 26,66. But given the above values, what are the percentages of the points in between 16 and 23 when 16 = 60% and 23 = 100%?

Answer by rothauserc(4718)

There are 7 increments between 16 and 23 which are to cover 40%
:
each point is 40 / 7 = 5.7143%
:
Note this approach assumes a uniform distribution of the 40%
:
so the 17 point is 65.7143%
:
looking at this as an arithmetic progression, then
:
x(n) = a(0) + d(n-17), where n is the point number, d=5.7143 and a(0) = 65.7143
:
the formula becomes
:
x(n) = 65.7143 + 5.7143(n-17)
: